### v1

# Preserving Multilingual Quality While Tuning Query Encoder on English Only

Abstract
A dense passage retrieval system can serve as the initial stages of information retrieval, selecting the most relevant text passages for downstream tasks. In this work we conducted experiments with the goal of finding how much the quality of a multilingual retrieval could be degraded if the query part of a dual encoder is tuned on an English-only dataset (assuming scarcity of cross-lingual samples for the targeted domain or task). Specifically, starting with a high quality multilingual embedding model, we observe that an English-only tuning may not only preserve the original quality of the multilingual retrieval, but even improve it.
Preserving Multilingual Quality While Tuning Query Encoder on English Only
Oleg Vasilyev, Randy Sawaya, John Bohannon Primer Technologies Inc. San Francisco, California oleg,randy.sawaya,john@primer.ai
1 Introduction
Advances in neural NLP methods have resulted in high quality dense vector text representations Reimers and Gurevych (2019); Cer et al. (2018); Conneau et al. (2017). Such representations are often used at the initial stages of an information retrieval system, selecting the most relevant documents, ranked relative to the query Xiong et al. (2020); Zhan et al. (2020, 2021); Ren et al. (2021b). A dual encoder is successfully used to train the representations Karpukhin et al. (2020); Ren et al. (2021a); Qu et al. (2021); Hofstätter et al. (2021); Ni et al. (2022); Dong et al. (2022). A dual encoder dense passage retrieval system is efficient for two main reasons: (1) it allows using simple inner product of query and document representations, and (2) it allows modifying the query representation for a task or domain, while keeping the stored and precomputed (query-invariant) document representations intact.
If the representation was pretrained in a multilingual setting, tuning on English-only samples may be expected to degrade the multilingual qualities and there may not be enough cross-lingual samples for tuning on a specific domain or types of queries. A multilingual query generator may be employed to overcome a shortage of cross-lingual data Ren et al. (2022); Zhuang et al. (2023), but, in this work, we follow an arguably simpler strategy. In order to understand the effect of English-only tuning on multilingual qualities of a representation, and to assess a possible degradation, we consider a simple setup: A state of the art multilingual embedding model is taken as the starting point, and fine-tuned by English only samples as a query representation part of the dual encoder.
Our contribution:
  1. 1.
    We show that fine-tuning of the query encoder on an English-only dataset may not only preserve multilingual qualities of query-document embeddings matching, but even improve them.
  2. 2.
    We consider tuning with parts of query encoder frozen, and suggest a conjectural explanation of the observations.
2 Setup
2.1 Model
In what follows, we use a state-of-the-art multilingual model intfloat/multilingual-e5-small [^1] https://huggingface.co/intfloat/multilingual-e5-small Wang et al. (2024) which we will call here $E5$. The model provides embeddings of a practical small size of $384$. For most of the evaluations, we also consider results using sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2222https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 Reimers and Gurevych (2019), referred to as $L12$; this model provides embeddings of the same size $384$.
2.2 Datasets
Here we made use of the following datasets:
  1. 1.
    MSMARCO Triplets [^3] https://huggingface.co/datasets/sentence-transformers/embedding-training-data/blob/main/msmarco-triplets.jsonl.gz, which we split on training, validation and test parts.
  2. 2.
    ARXIV subset, which we take from a mirror of arxiv (version 173) [^4] https://huggingface.co/datasets/arxiv-community/arxiv_dataset [^5] https://www.kaggle.com/datasets/Cornell-University/arxiv
  3. 3.
    XNLI multilingual dataset [^6] https://huggingface.co/datasets/facebook/xnli Conneau et al. (2018), of which we use the test subset.
The MSMARCO dataset consists of 499184 samples, with each sample being a tuple given as (query, positives, negatives). The “positives” are the correct answers to the query, and the “negatives” are semantically similar, but incorrect answers. For most samples there is only one positive, but many negatives. For our tuning we simply select the very first positive and the very first negative. Thus, each sample gives one triplet (anchor, positive, negative) for contrastive learning, where the query is taken as an anchor. We keep the first 487983 samples (or 34856 batches if each batch is 14 triplets) for tuning, leaving the next 4200 samples (300 batches) for validation, and the last 7000 samples for evaluation. During evaluation we create all possible triplets from the 7000 samples (using all positives and negatives); this makes 357642 evaluation triplets.
Of ARXIV we use titles and abstracts. In order to have a representative subset of a manageable size for our evaluations, we select all samples that have at least one category with a maximum size of 10K samples. For example, the arxiv category bayes-an is the smallest (size 16) in our snapshot (version 173), meaning that there were only 16 arxiv preprints in this category.
We made two flavors of evaluation arxiv triplets from this arxiv subset. In the first version, the anchor is the title, the positive is the corresponding abstract, and the negative is another random abstract. In the second version the anchor is the first sentence of an abstract, the positive is the rest of the abstract, and the negative is a similar piece (first sentence excluded) of another, random (with fixed seed) abstract. (For full reproducibility of obtaining arxiv triplets, see Appendix A.)
In evaluating on triplets (of MSMARCO and ARXIV) we compare the similarity (or distance) between the anchor and the positive vs the negative. If the positive does not turn out to be closer than the negative to the anchor, we count this as an error.
XNLI consists of pairs of sentences which are human-labeled as entailment, neutral or contradiction. The test subset (which we use) contains 1670 pairs for each of these labels and each sentence is presented in 15 languages: [’ar’, ’bg’, ’de’, ’el’, ’en’, ’es’, ’fr’, ’hi’, ’ru’, ’sw’, ’th’, ’tr’, ’ur’, ’vi’, ’zh’]. We use 225 versions of the pairs, because each sentence of the pair can be in any of the 15 languages. At evaluation the first sentence serves as the query (the embedding is taken by the query model), and the second one as the text. We expect that the sentences of an entailment pair should be closer to each other than the sentences of any neutral pair, or of any contradiction pair. Whenever this does not happen, we count this as an error.
data Evaluation $E5$ $L12$
MM N tot 357642
N errors (cos) 16688 54156
N errors (dist) 17036 55103
ARX-F N tot 253140
N errors (cos) 4014 12488
N errors (dist) 3939 17009
ARX-T N tot 253140
N errors (cos) 399 3547
N errors (dist) 453 4264
XNLI N total 2788900
entail-neutral (cos) 10.8% 10.2%
entail-contr (cos) 10.0% 7.2%
entail-neutral (dist) 10.5% 10.1%
entail-contr (dist) 9.6% 7.8%
Table 1: The count of errors for the original untuned models $E5$ and $L12$, on the datasets noted in the first column: MM - MSMARCO last 7000 samples (357642 triplets, see Section 2.2); ARX-F - arxiv subset with the abstract’s first sentence as an anchor; ARX-T - arxiv subset with the title as an anchor; XNLI - XNLI test subset providing 1670x1670=2788900 comparisons of entailment pairs vs neutral pairs (and the same amount of entailment pairs vs contradiction pairs). For XNLI the errors are averaged over 225 (15x15) language-language versions, and shown as percent of $Ntotal$.
2.3 Tuning
We tune a dual encoder by contrastive learning on MSMARCO in the following simple regime:
  1. 1.
    The text encoder is fully frozen; some parts of the query encoder can be frozen.
  2. 2.
    Unless specified otherwise, the batch size is 14, the learning rate is 5e-8 and the contrastive learning margin is $0.1$. The loss is defined by the triple margin loss.
  3. 3.
    Unless specified otherwise, there are $1000$ batches per epoch, and stopping occurs after $10$ consecutive non-improvement epochs. The improvement is measured on the validation subset after each epoch. The model is considered to be improved if (on the validation subset) both the loss and the count of errors are decreased.
We have extensively considered changing to this tuning parameters in Appendix  B (and in Appendix  C for model $L12$); variations of tuning scheduling are considered in Appendix  D.
3 Observations
3.1 Original multilingual embedding
To establish a baseline before any fine-tuning, and to ensure our evaluation is not too easy, we measure the errors of the original $E5$ model on the data described in Section 2.2 and show the results in Table 1. (We also measure the errors of $L12$, which does far worse on English data (MSMARCO, ARXIV) by comparison.)
The count of errors on the triplets (MSMARCO, ARXIV) is straightforward: it is an error when a positive is not closer than a negative to the anchor of the triplet. On XNLI we sum up the error count over all language-language pairs and divide the sum by the number (255=15x15) of such pairs. This averaged error is shown as a percentage of the total (2788900) comparisons; each comparison here is either a comparison of an entailment-labeled sample with a neutral-labeled sample (entail-neutral in the table) or a comparison of an entailment-labeled sample with a contradiction-labeled sample (entail-contr in the table). An error was counted whenever the sentences of an entailment sample happened to be farther from each other than the sentences of a neutral (or contradiction) sample.
The amount of errors (Table 1) is reasonable enough to consider how tuning would affect them.
msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
frozen cdcdcdc+/- d+/- c+/- d+/-
- 7.47 8.46 4.51 4.85 1.75 5.52 222/0 215/2 194/4 147/23
emb.base 7.32 8.82 4.16 5.08 3.51 7.73 222/0 217/1 201/2 159/21
emb 7.30 8.82 4.21 5.05 3.51 7.73 222/0 217/1 201/2 158/20
emb, B0a 7.30 8.76 4.09 5.00 3.26 7.73 222/0 217/1 200/2 159/21
emb, B0a,i 7.48 9.00 4.36 5.03 3.26 7.73 223/0 219/0 199/2 156/25
emb, B0a,i,od 7.31 8.82 4.09 4.85 3.51 7.73 222/0 217/1 200/2 158/21
emb, B0 7.35 8.78 4.09 5.10 3.51 7.73 222/0 217/1 200/2 159/19
emb, B0-5 7.87 9.39 5.11 5.74 3.26 7.51 219/0 213/3 200/5 157/25
emb, B0-10 1.45 2.57 0.17 0.86 0.00 0.44 123/0 112/0 21/0 25/10
Table 2: Evaluations of the $E5$ query model tuned on MSMARCO as described in Section 2.3. The rows are in the order of increased freezing (at tuning): from no freezing (top row) to freezing everything up to the last transformer block (the last row); the last block is $B11$. The $emb.base$ model has only the first three layers of the embedding block frozen (tokens, positions, token-types). The $emb$ model has the full embedding block frozen. The other notations: $B0$ is the full first transformer block; $B0$-$5$ are the first 6 blocks; the extensions $a$, $i$, $od$ (for $B0$) correspond to the layers $attention$, $intermediate$ and $output.dense$ of the block. The columns $c\%$ and $d\%$ show the improvement (in percents) relative to the original $E5$ model by evaluation with cosine ($c$) and distance ($d$). The columns $c$+/- and $d$+/- show how many language pairs are improved ($+$) or worsened (--) by statistically significant change of the number of errors, measured with cosine ($c$) or distance ($d$). Values that did not pass two-tailed test are gray-colored.
3.2 Tuning partially frozen query model
Tuning (as specified in Section 2.3) of the dual encoder with the text encoder frozen and query model free ends up with the result shown in the first row of Table 2. The improvement $I$ of the performance on triplets (columns $c\%$ and $d\%$ in the table) is presented by the ratio
$I=\frac{N_{e}-\tilde{N_{e}}}{N_{e}}$ (1)
where $N_{e}$ is the original number of errors, and $\tilde{N_{e}}$ is the number of errors after the tuning. The improvement $I$ is shown as a percentage and is measured either by cosine (columns $c\%$) or by distance (columns $d\%$).
The improvement on XNLI is shown by the number of language pairs in which the number of errors changed in a statistically significant way. In the columns c+/- (cosine similarity is used) and d+/- (distance is used) the number of improved language pairs is shown in the “numerator” (+), and the number of worsen pairs is shown in the “denominator” (-) in the cell. The change for a language-language pair is considered statistically significant if the Z-value for two-tailed test exceeds $1.96$ (which corresponds to 0.025 on each tail).
For example, the first row of Table 2 shows that the tuning of the query model with no frozen parts improves its cosine-measured performance on the MSMARCO evaluation samples by $7.47\%$; it also improves on ARXIV, and, interestingly, it has shown more improvement than worsening on multilingual XNLI data. For cosine-assessed comparisons of entailment-labeled and neutral-labeled XNLI samples, the tuning statistically significantly improved 222 out of 225 language pairs, without worsening a single one. Of the XNLI columns the worst result is in the last one: the distance-assessed comparisons of entailment-labeled and contradiction-labeled pairs. It still shows more improved than worsened language pairs (147 vs 23).
In the next rows of Table 2 we increase the frozen part of the query model, guided by the intuition that the lowest, closest to input, layers (especially the embedding block) may be dealing with specifics of the languages, while the higher transformer blocks may be less dependent on the languages. Indeed, as we freeze the lower three layers of the embedding block (word, position and token type embedding layers), we see even better improvement on XNLI. The third row $emb$ corresponds to freezing the whole embedding block i.e. freezing its $LayerNorm$ in addition to $emb.base$); this gives similar results.
Freezing beyond the embedding block in the query model, we freeze the embedding block together with parts of the first transformer block, such as its attention layers ($B0_{a}$), then also its intermediate layers ($B0_{a,i}$), then also output dense layers ($B0_{a,i,od}$), and finally the whole block $B0$. The XNLI improvement remains about as good as when only the embedding block was frozen ($emb$ row). Freezing half of the transformer ($emb$ and $B0$-$5$) still gives similar results, if only slightly worse on XNLI. However, freezing everything up to the last transformer block (last row in Table  2) naturally diminishes the overall effect of the tuning. For verifying our intuition, we also tried the opposite to freezing the embedding: we had frozen everything except the embedding block. In such setting our tuning regime could not improve the model at all.
In Appendix E we show improvements for each XNLI language pair, for the case of our default tuning (Section 2.3), with embedding block frozen.
Our results do not depend strongly on the tuning parameters (Appendix B), unless the parameters are changed too much, especially the contrastive learning triple loss margin (Appendix B.3). Naturally, increasing learning rate delivers more gains on MSMARCO and ARXIV, while slightly reducing gains on XNLI. The learning rates between 2e-8 and 6e-8 may be considered as the best balance, but a rate as high as 1e-7 may be preferred, depending on languages of interest, see Appendix B.1.
The batch sizes as different as 7, 14, 28, 56 and 112 deliver similar gains, if the number of samples per epoch 14000 is kept the same (Appendix B.2, Table 4). The batch size 112 is ’safer’, giving less gains but not degrading any language pairs if evaluated with cosine similarity, and only single one when evaluated with distance. If the number of batches 1000 (rather than number of samples) is kept the same (Table 5 then the large batch size 112 delivers highest gain on MSMARCO, but far worse results on ARXIV and XNLI.
The margin of contrastive learning is good between 0.05 and 0.15 (Appendix B.3). Increasing it beyond 0.15 diminishes gains in all the datasets, and at margin 0.5 the number of degraded language pairs becomes comparable to the number of improved language pairs.
Varying our stopping criterion (10 consecutive non-improvement epochs) between 5 and 15 does not have much effect (Appendix B.4). Interestingly, switching to a less straightforward tuning regime, i.e. using scheduling and weight decay may deteriorate the performance gains on MSMARCO and ARXIV, and may degrade performance on many language pairs in XNLI.
The model $L12$, having much more errors on English samples at start (Table 1), also can improve/preserve performance on XNLI, but not as good as $E5$ (Appendix C). We speculate that this may be related to its worse starting point.
4 Conclusion
We considered tuning the query part of a dual encoder starting from a high quality multilingual embedding model, and using English-only samples in the tuning. We found that multilingual qualities are quite stable in many scenarios of the tuning, and can be not only preserved but improved. We explain this by speculating that most of the transformer, except the embedding block, depends weakly on multiple languages. We also think of this as a particular case of a more general pattern: tuning a certain model quality, if done carefully enough, can improve the related qualities too. In our setting this allows the simple and resource-light adjustment of multilingual embeddings for specific query types or domains.
5 Limitations
Our consideration here is limited to starting with a single high quality multilingual embedding model, and tuning it (on English-only samples) as a query encoder. While this setup is good for understanding and convenient for adjusting an existing model, it would be natural to follow this up by considering a pre-trained multilingual dual encoder which is already asymmetric from the start.
For our illustration we used the state of the art multilingual model intfloat/multilingual-e5-small, and also, for comparison, repeated the same observations for the sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 model. To gain a better understanding of the observed behaviours, it would be good to investigate more multilingual models.
We considered tuning the query encoder on English-only samples, and found that such tuning can “pull up” the quality of other languages too. Choosing another language for tuning would be interesting both for understanding and as a practical scenario.
We used MSMARCO triplets for tuning; for evaluation we used a set aside part of MSMARCO triplets, and ARXIV in two variations, and XNLI. The motivation was that the MSMARCO evaluation part must show improvement (after tuning), ARXIV must verify the robustness of the improvement on a very different kind of texts, and XNLI must reveal the effect of the English-only driven improvement on multilingual qualities. That said, the evaluations can be extended to more datasets.
Appendix A Triplets from ARXIV subset
We make use of triplets created from arxiv because this provides our evaluation with a very different kind of text (compared to MSMARCO), and thus allows us to judge robustness of the improvement. The procedure of making the triplets is simple, and it may be more appropriate to provide a recipe rather than duplicate already existing data: arxiv-metadata-oai-snapshot.jsonl from mirror of arxiv (version 173). As outlined in Section 2.2, the triplets are created in the following steps:
  1. 1.
    Identify all arxiv categories with a maximum size of 10K articles (i.e. arxiv preprints).
  2. 2.
    Select in the same order as given all samples (articles) that have at least one of the categories identified above. This is the subset of arxiv to deal with: manageably small, yet diverse. Keep the order of the samples in the subset, and iterate over them in this order as following:
  3. 3.
    For each sample $A$ in the subset, pick a “random” sample $B$ (random seed 0, python version 3.8). If the picked up $B$ coincides with $A$, select instead the sample next after $B$. If no next, select the first sample. Create a triplet either as (1) the title of sample $A$ as the anchor, its abstract as positive, and abstract of sample $B$ as negative, or (2) the first sentence of the abstract of sample $A$ as the anchor, the rest of the abstract as positive, and the abstract of sample $B$ (first sentence excluded) as negative.
The difference made by the “randomness” of the selection of negatives is negligible, but if exact reproducibility on arxiv is wanted, fixing the random seed to 0 in python 3.8 should do it.
learning msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
rate cdcdcdc+/- d+/- c+/- d+/-
1e-8 0.73 0.94 -0.35 0.00 0.50 -0.44 0/0 0/0 0/0 0/0
2e-8 6.50 7.62 4.71 4.75 2.76 5.96 222/0 218/0 208/1 171/5
3e-8 6.65 7.57 4.56 4.57 3.51 5.96 222/0 217/0 202/1 170/6
4e-8 6.66 7.58 4.31 4.62 2.76 6.18 222/0 218/0 207/1 174/5
5e-8 7.30 8.82 4.21 5.05 3.51 7.73 222/0 217/1 201/2 158/20
6e-8 7.14 8.74 4.63 5.10 3.26 7.28 222/0 218/0 204/2 168/13
7e-8 9.11 11.35 5.58 5.84 3.01 8.83 223/0 216/3 193/17 140/52
8e-8 8.75 10.98 5.75 6.50 4.51 9.05 222/0 217/3 186/20 131/61
9e-8 9.01 11.26 5.78 6.17 4.01 8.83 222/0 216/3 182/25 129/67
1e-7 9.26 11.38 5.38 6.12 3.51 9.49 222/0 214/3 188/18 132/63
Table 3: Evaluations of the $E5$ query encoder tuned with a frozen embedding block, batch size 14, margin 0.1 and different learning rates (first column). Values that did not pass the two-tailed test are shown in gray.
batch msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
size cdcdcdc+/- d+/- c+/- d+/-
7 6.85 7.57 3.86 4.44 3.01 6.62 221/0 217/0 198/2 163/13
14 7.30 8.82 4.21 5.05 3.51 7.73 222/0 217/1 201/2 158/20
28 7.45 9.47 4.46 5.13 4.01 8.39 222/0 219/1 196/6 145/27
56 6.51 7.33 3.46 4.14 2.51 5.74 221/0 217/0 202/1 168/6
112 4.63 4.78 1.84 2.16 2.51 2.21 212/0 203/0 196/0 155/1
Table 4: Evaluations of the $E5$ query encoder tuned with a frozen embedding block, learning rate 5e-8, margin 0.1 and different batch sizes (first column); 14000 samples per epoch. Values that did not pass the two-tailed test are shown in gray.
batch msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
size cdcdcdc+/- d+/- c+/- d+/-
7 6.76 7.99 3.69 4.72 3.26 6.18 221/0 216/0 177/7 119/31
14 7.30 8.82 4.21 5.05 3.51 7.73 222/0 217/1 201/2 158/20
28 8.50 10.18 5.03 6.12 2.76 8.17 222/0 218/3 197/9 147/36
56 7.42 9.12 3.86 4.47 3.76 8.17 223/0 220/0 200/2 160/21
112 9.50 11.84 -1.54 -4.42 -15.54 -14.13 175/35 146/65 91/117 32/175
Table 5: Evaluations of the $E5$ query encoder tuned with a frozen embedding block, learning rate 5e-8, margin 0.1 and different batch sizes (first column); 1000 batches per epoch. Values that did not pass the two-tailed test are shown in gray.
margin msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
cdcdcdc+/- d+/- c+/- d+/-
0.05 7.40 9.32 3.71 3.83 2.76 5.30 225/0 223/0 201/4 169/21
0.10 7.38 8.95 4.91 4.90 3.26 7.73 222/0 218/0 201/2 163/17
0.15 6.90 7.78 6.00 5.79 4.26 7.51 213/0 204/3 176/7 125/33
0.20 6.40 7.23 7.32 6.80 5.01 8.39 175/6 161/14 132/32 94/67
0.25 5.09 5.70 7.42 7.11 5.51 7.73 127/27 116/42 108/55 76/84
0.30 3.92 4.23 6.50 6.35 4.51 6.40 108/40 99/52 97/62 79/85
0.35 3.76 3.80 4.53 4.54 4.01 3.53 80/40 74/49 72/57 65/75
0.40 2.77 3.10 5.06 5.41 4.51 5.30 69/72 67/69 64/80 61/94
0.45 2.59 2.78 4.68 4.85 4.51 4.64 54/81 59/77 58/83 56/98
0.50 2.94 2.41 2.99 2.44 2.51 1.77 30/23 31/27 53/19 48/26
Table 6: Evaluations of the $E5$ query encoder tuned with a frozen embedding block, learning rate 5e-8, batch size 14 and different triplet loss margins (first column). Values that did not pass the two-tailed test are shown in gray.
idle epochs msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
to stop cdcdcdc+/- d+/- c+/- d+/-
5 6.50 7.60 4.09 4.60 2.26 6.62 222/0 217/0 208/1 174/5
10 7.38 8.95 4.91 4.90 3.26 7.73 222/0 218/0 201/2 163/17
15 8.93 10.98 5.83 6.07 2.51 7.73 222/0 217/3 191/17 140/51
Table 7: Evaluations of the $E5$ query encoder tuned with a frozen embedding block, learning rate 5e-8, batch size 14 and triplet loss margin 0.1, stopped after different number of idle epochs (first column). The epoch is idle if no improvement is made. Values that did not pass the two-tailed test are shown in gray.
msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
frozen cdcdcdc+/- d+/- c+/- d+/-
- 6.60 6.93 2.65 -7.86 14.46 -0.12 206/15 57/35 201/15 15/189
emb.base 7.28 8.04 2.46 -9.17 12.38 -1.03 200/15 47/47 206/15 20/142
emb 7.28 8.04 2.51 -9.17 12.4 -1.03 200/15 47/47 206/15 20/143
emb, B0a 7.26 8.03 2.29 -9.22 12.52 -0.96 201/15 46/46 206/15 20/142
emb, B0a,i 7.03 7.75 2.44 -8.6 12.46 -0.63 203/15 51/41 206/15 20/133
emb, B0a,i,od 9.04 10.15 1.80 -17.54 12.49 -8.58 195/19 30/116 207/15 19/167
emb, B0 8.92 9.98 1.78 -16.73 12.88 -8.07 195/16 33/102 209/15 20/163
emb, B0-5 8.54 9.68 2.71 -19.01 12.35 -12.17 209/15 28/129 209/15 19/172
emb, B0-10 0.11 0.15 0.10 -0.12 0.25 -0.02 0/0 0/0 0/0 0/0
Table 8: Evaluations of the $L12$ query model tuned on MSMARCO as described in Section 2.3. The rows are in the order of increased freezing (at tuning): from no freezing (top row) to freezing everything up to the last transformer block (the last row); the last block is $B11$. The $emb.base$ model has only the first three layers of the embedding block frozen (tokens, positions, token-types). The $emb$ model has the full embedding block frozen. The other notations: $B0$ is the full first transformer block; $B0$-$5$ are the first 6 blocks; the extensions $a$, $i$, $od$ (for $B0$) correspond to the layers $attention$, $intermediate$ and $output.dense$ of the block. The columns $c\%$ and $d\%$ show the improvement (in percents) relative to the original $L12$ model by evaluation with cosine ($c$) and distance ($d$). The columns $c$+/- and $d$+/- show how many language pairs are improved ($+$) or worsened (--) by statistically significant change of the number of errors, measured with cosine ($c$) or distance ($d$). Values that did not pass two-tailed test are gray-colored.
learning msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
rate cdcdcdc+/- d+/- c+/- d+/-
1e-8 1.85 2.15 0.76 -1.17 3.75 0.63 55/15 0/15 131/15 0/16
2e-8 4.81 5.18 1.77 -3.98 8.32 1.22 199/15 41/25 196/15 6/59
3e-8 4.70 5.11 1.94 -4.03 8.74 1.29 198/15 39/25 195/15 6/64
4e-8 7.22 8.03 2.23 -8.98 12.21 -1.13 203/15 50/41 207/15 20/134
5e-8 7.28 8.04 2.28 -9.23 12.4 -1.03 200/15 47/47 206/15 20/143
6e-8 7.16 7.91 2.55 -9.07 12.69 -0.98 203/15 52/41 207/15 20/135
7e-8 9.41 10.72 0.88 -20.77 10.94 -12.1 174/28 20/174 201/16 16/185
8e-8 9.45 10.69 1.23 -20.97 11.64 -12.03 178/28 25/164 201/15 16/182
9e-8 9.29 10.57 2.03 -19.69 12.88 -9.92 187/25 28/144 203/15 17/178
1e-7 9.40 10.74 1.84 -20.44 12.43 -10.69 180/27 28/160 202/15 17/177
Table 9: Evaluations of the $L12$ query encoder tuned with a frozen embedding block, batch size 14, margin 0.1 and different learning rates (first column). Values that did not pass the two-tailed test are shown in gray.
batch msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
size cdcdcdc+/- d+/- c+/- d+/-
7 8.74 9.89 1.59 -16.43 9.59 -9.76 195/18 33/106 205/16 18/168
14 7.28 8.04 2.28 -9.23 12.4 -1.03 200/15 47/47 206/15 20/143
28 5.02 5.49 1.98 -4.05 8.37 1.27 199/15 36/27 196/15 7/77
56 5.05 5.35 1.73 -4.3 8.66 0.63 197/15 35/28 195/15 6/89
112 4.68 5.01 1.69 -3.64 7.47 0.77 193/15 25/25 188/15 4/86
Table 10: Evaluations of the $L12$ query encoder tuned with a frozen embedding block, learning rate 5e-8, margin 0.1 and different batch sizes (first column); 14000 samples per epoch. Values that did not pass the two-tailed test are shown in gray.
batch msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
size cdcdcdc+/- d+/- c+/- d+/-
7 6.47 8.48 -1.79 -24.53 -1.1 -25.33 12/198 3/212 147/52 9/208
14 7.28 8.04 2.28 -9.23 12.4 -1.03 200/15 47/47 206/15 20/143
28 9.54 10.78 1.11 -20.67 9.16 -13.06 172/29 18/175 201/16 16/185
56 9.51 10.86 1.04 -21.01 8.37 -13.63 173/28 18/175 201/16 17/181
112 9.44 10.88 1.06 -20.86 8.18 -13.77 174/28 21/174 200/17 15/188
Table 11: Evaluations of the $L12$ query encoder tuned with a frozen embedding block, learning rate 5e-8, margin 0.1 and different batch sizes (first column); 1000 batches per epoch. Values that did not pass the two-tailed test are shown in gray.
margin msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
cdcdcdc+/- d+/- c+/- d+/-
0.05 9.00 10.16 1.09 -22.22 10.23 -13.81 180/28 27/159 205/15 21/171
0.10 7.28 8.04 2.28 -9.23 12.4 -1.03 200/15 47/47 206/15 20/143
0.15 9.90 11.34 1.11 -19.64 11.98 -10.18 164/29 19/177 193/20 15/195
0.20 9.71 11.16 1.58 -16.94 12.63 -6.99 163/29 20/171 193/22 15/198
0.25 9.30 10.75 1.39 -13.92 12.83 -4.2 163/27 16/160 183/22 14/205
0.30 9.62 11.1 1.21 -13.82 12.6 -3.94 125/33 13/180 157/31 14/209
Table 12: Evaluations of the $L12$ query encoder tuned with a frozen embedding block, learning rate 5e-8, batch size 14 and different triplet loss margins (first column). Values that did not pass the two-tailed test are shown in gray.
msmarco arxiv-first arxiv-title xnli ent-neutr xnli ent-contr
B Sch D cdcdcdc+/- d+/- c+/- d+/-
100 $Q$1.12 -1.17 -0.82 -0.14 -0.52 -1.71 0/138 3/117 15/65 61/40
$E_{0.98}$ $10^{-6}$0.31 -0.18 -0.39 -0.05 -0.52 -1.22 0/0 0/0 0/11 0/7
64 - - -1.20 -1.05 -0.79 -0.11 -0.26 -0.73 0/57 5/34 1/80 9/47
$Q$0.72 -0.89 -0.66 -0.14 -0.26 -1.71 0/65 4/43 1/62 21/33
$E_{0.98}$ $10^{-6}$0.96 -1.13 -0.66 -0.49 -0.52 -1.22 0/80 6/56 3/65 29/35
$E_{0.95}$ $10^{-6}$1.78 -2.36 -0.97 -1.35 0.78 -1.46 1/160 9/131 22/103 73/68
$E_{0.95}$ $10^{-5}$1.78 -2.36 -0.97 -1.35 0.78 -1.46 1/160 9/131 22/103 73/68
- $10^{-4}$1.20 -1.05 -0.79 -0.11 -0.26 -0.73 0/57 5/34 1/80 9/47
$L$ $10^{-4}$0.76 -0.80 -0.58 -0.16 -0.26 -1.46 0/61 3/32 0/58 14/32
32 $Q$ $10^{-4}$2.12 -3.35 -1.47 -1.60 0.00 -2.68 2/190 8/162 41/107 93/69
$L$ $10^{-4}$2.05 -3.36 -1.45 -1.33 0.00 -2.68 2/189 8/158 41/102 94/68
$E_{0.98}$ $10^{-4}$2.12 -3.54 -1.42 -1.84 0.00 -2.68 2/192 8/163 41/110 93/71
16 $L$0.02 0.13 -0.58 -0.65 -1.04 -1.22 0/0 0/0 11/0 43/0
$E_{0.95}$2.52 -3.32 -1.32 -1.38 -0.26 -2.20 3/186 8/164 49/86 101/55
$E_{0.95}$ $10^{-4}$2.52 -3.32 -1.32 -1.38 -0.26 -2.20 3/186 8/164 49/86 101/55
8 $Q$ $10^{-6}$2.05 -3.29 -1.11 -1.38 -0.26 -3.66 0/212 3/182 21/140 70/105
$E_{0.95}$ $10^{-6}$2.44 -3.81 -1.55 -1.78 -0.52 -3.41 0/213 4/182 24/135 76/97
- $10^{-5}$2.03 -3.19 -0.74 -1.57 -0.26 -3.41 0/209 3/182 20/139 71/103
$Q$ $10^{-4}$2.05 -3.29 -1.11 -1.38 -0.26 -3.66 0/212 3/182 21/140 70/105
$E_{0.95}$ $10^{-4}$2.44 -3.81 -1.55 -1.78 -0.52 -3.41 0/213 4/182 24/135 76/97
Table 13: Percentage improvement over the fine-tuned E5 model with a frozen embedding block and tuned using a batch size of $14$, learning rate $10^{-7}$ and a margin of $0.1$. The blue colors indicate the top improvements whereas the red colors indicate the worse degradation. Three parameters are randomly varied: the batch size (denoted as “B”), the learning rate scheduler (denoted as “Sch”) and the weight decay (denoted as “D”). The learning rate schedulers are defined in Table 14 with an initial learning rate of $10^{-7}$. cand drefer to measuring the similarity of the text pairs using either the cosine similarity or the euclidean distance, respectively. For XNLI, (+)(+)( + ) indicates the number of language pairs that were improved while ()(-)( - ) indicates those that have worsened out of a total of 225 language pairs. Note that only the statistically significant (determined by a Z-test) language pairs are retained and hence not all the improved/worsened counts sum to 225. Additionally, (ent-neutr) refers to entailment-entailment similarities compared with entailment-neutral similarities whereas (ent-contr) refers to comparisons against entailment-contradiction similarities.
Appendix B Dependency on tuning parameters
B.1 Learning rate
Our results (Section 3, Table 2) do not depend strongly on the tuning parameters, but an order of magnitude change in learning rate does have some effect. As shown in Table 3, the rates in the range 2e-8 to 6e-8 provide similar results; the lower rates 2e-8 to 4e-8 may be slightly safer on languages, with weaker gain on English data. The rate 1e-8 is too low: tuning barely made any changes. The higher rates 7e-8 to 1e-7 give higher gains on English data, but the number of degraded language pairs for cosine matching jumps from 1-2 to around 20-25 (out of 225).
B.2 Batch size
In Table 4 we show results for batch sizes 7, 14, 28, 56 and 112, while keeping the number of samples per epoch the same (14000). The row with batch 14 here coincides with row with learning rate 5e-8 in Table 3, and with row for frozen embedding block in Table 2. The results for all batch sizes are similar. Tuning with the higher batch size 112 is a bit ’safer’ for languages, not degrading any language pair when evaluated by cosine measure, and degrading only one language pair (for entailment vs contradiction) when evaluated by distance measure. This comes at the price of lower gains on MSMARCO and ARXIV.
Table 5 shows what happens if the number of batches per epoch (1000) is kept the same, rather than the number of samples. In this setting the larger batch size 112 leads at tuning to a less frequent validation (by MSMARCO validation subset) and, effectively, to later and less reasonable stopping. This results in higher gain on MSMARCO test subset, but in far worse results on ARXIV and XNLI.
B.3 Margin of triple loss
The margin is an important parameter of the contrastive learning through the triplet loss. In Table 6 we show the dependency of the evaluation results on the margin during its tuning. We consider our default tuning parameters (Section 2.3), but change the margin. The results are natural: a margin up to 0.15 is reasonable, and at higher margins the disturbance on cross-lingual, and, eventually, on English data evaluation gets too strong.
B.4 Stopping criterion
In Table 7 we show dependency of improvement results on the stopping criterion. The stoppings after 5 or 10 non-improvement epochs give similar results. Stopping after 15 non-improvement epochs continues the trend of increased gain on English data, but with a deterioration for a few language pairs.
Appendix C L12 Model
For comparison with our consideration (Section 3) of the $E5$ model, here we show a few results of the query model tuning (in the same regime, see Section 2.3) that instead starts with the $L12$ model. Altogether, we observe that $L12$ does not preserve languages as nicely as $E5$, even for the cosine measure. And yet the tuning of $L12$ makes much more language pairs improved than degraded. While each language pair is important in its own right, we find it interesting that even for $L12$ the change in the absolute number of errors summed up across all language pairs is still within $1\%$.
C.1 Freezing layers
Table 8 shows results of tuning with freezing some of $L12$ layers. It is similar to the Table 2 for $E5$. And, similar to $E5$, freezing everything except the embedding, resulted in negligible changes of the query encoder (not shown in the table).
C.2 Learning rate
In Table 9 we show evaluations of $L12$ model tuned at different learning rates. This table is arranged similarly to the $E5$ model Table 3. The results for $L12$ are not as good as for $E5$, this is especially noticeable by the number of language pairs that have degraded.
C.3 Batch size
Dependency of $L12$ tuning on the batch size is shown in Table 10 (number of samples per epoch is 14000) and in Table 11 (number of batches per epoch is 1000). Observations are somewhat similar to $E5$ (Appendix B.2, except that generally $L12$ performs not as good as $E5$ and the batch size 7 turns out bad for $L12$.
C.4 Margin of triple loss
Dependency of $L12$ tuning on the contrastive learning margin size is shown in Table 12. The results for margin 0.1 are distinctly better. Altogether, the model $L12$ appears to be more sensitive (compared to $E5$) to the tuning parameters if the goal is to preserve performance on multilingual data XNLI and on out-of-domain data ARXIV.
Appendix D Effects of Learning Rate Scheduler and Weight Decay
Using the fine-tuned E5 model with the frozen embedding block, tuned using a batch size of 14, learning rate of $10^{-7}$ and a margin of 0.1, we randomly vary the batch size, learning rate scheduler and weight decay in order to assess their impact on the model’s final performance. In Table 13, we present the change in performance across these different configurations. Table 14 lists the different schedulers we considered. Values in blue indicate the top improvements whereas values in red indicate the worse degredations.
Across these parameters, we see that, on average, the batch size appears to have to most significant impact, generally leading to poorer performance as the batch size is decreased. Within each batch size group, we see that using an exponential learning rate scheduler($E_{0.95}$ or $E_{0.98}$) is generally worse than using any of the other schedulers or no scheduler at all. A specific exception exists when using a batch size of 100 where the exponential scheduler outperforms the quadratic one.
Scheduler Definition
$L$ $\alpha(t)=\alpha_{0}\left(1-\frac{t}{T}\right)$
$Q$ $\alpha(t)=\alpha_{0}\left(1-\left(\frac{t}{T}\right)^{2}\right)$
$E_{0.95}$ $\alpha(t)=0.95^{t}\alpha_{0}$
$E_{0.98}$ $\alpha(t)=0.98^{t}\alpha_{0}$
Table 14: The definitions of the various learning rate schedulers used in Table 13 where $t$ is the current training step, $T$, the total number of training steps and $\alpha_{0}$, the initial learning rate.
Appendix E Improvement on XNLI, by language pairs
In Table 2 and in other similar tables we have been showing the improvement on XNLI dataset by a number of improved or degraded language pairs, when the changes are statistically significant.
Figure 1: Improvement of $E5$ on XNLI entailment-neutral comparisons assessed by cosine, shown by language pairs. The improvement results from tuning on MSMARCO with embedding block frozen and tuning parameters as specified in Section 2.3.
Figure 2: Improvement of $E5$ on XNLI entailment-contradiction comparisons assessed by cosine.
Figure 3: Improvement of $L12$ on XNLI entailment-neutral comparisons assessed by cosine.
Figure 4: Improvement of $L12$ on XNLI entailment-contradiction comparisons assessed by cosine.
Here we have shown in Figures 123 and  4 the actual improvement calculated for each language pair separately, by equation 1, using the count of errors for the language pair on XNLI.